OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 25 feet and the child ride has a radius of 10 feet, what is the ratio between the volumes of each ride? options: 25:4 15:6 125:8 5:2

OpenStudy (anonymous):

please help

hero (hero):

Just use the formula for volume of a cone to find the ratio: \[\frac{\text{Volume of Cone One}}{\text{Volume of Cone Two}} = \frac{\pi r^2h/3}{\pi r^2h/3} = \frac{\pi (25)^2h/3}{\pi(10)^2h/3}\] Basically, everything cancels except \[\frac{25^2}{10^2}\] If you simplify that, you might get one of your answer choices.

OpenStudy (anonymous):

thanks @Hero

hero (hero):

Did you figure out which answer choice is correct?

OpenStudy (anonymous):

wait i got 6.25 and thats not any of my choices @Hero

hero (hero):

That's because you simplified it to a decimal form rather than fraction form.

OpenStudy (jdoe0001):

rationalize 6.25

OpenStudy (anonymous):

so wat would it be in fraction?

OpenStudy (jdoe0001):

hmm, so you're doing rations and dunno how to simplify fractions?

OpenStudy (anonymous):

no thats y i need help

hero (hero):

\[\frac{25^2}{10^2} = \left(\frac{25}{10}\right)^2 = \left(\frac{5 \times 5}{5 \times 2}\right)^2 = \left(\frac{5}{2}\right)^2 = \frac{25}{4}\]

OpenStudy (anonymous):

ohh i get it now thanks a lot